# Evolving Generalizable Parallel Algorithm Portfolios via Domain-Agnostic Instance Generation

### v3

@@ 15             $\Psi\leftarrow\Psi\cup\left\{\theta^{\prime}\right\}$;
20       Assign the fitness of each $m\in M$ as $-f(P,m)$;
@@ 22       $M_{new}\leftarrow\emptyset$;
25             Test $P$ on $m^{\prime}$ and assign the fitness of $m^{\prime}$ as $-f(P,m^{\prime})$;
@@
28             $M\leftarrow M\backslash\left\{m^{\star}\right\}$;
@@ III-C NIR-based Instance Generation and Evaluation
Since the performance indicator $f$ typically lacks analytic forms and $\boldsymbol{e}^{\prime}$ is a real-valued vector, Eq. (10) represents a continuous black-box optimization problem. In this work, PGPE [37], an evolution strategy (ES) method, is employed to optimize Eq. (10) (lines 1-2 in Alg. 1). Details of PGPE are provided in Appendix A of the supplementary. Note that other black-box continuous optimizers could also be used here, as the specific choice of optimizer is not central to our approach. Upon termination, the mutation operator returns the NIR $m^{\prime}$ specified by the best embedding $\boldsymbol{e}^{\prime}$ found by the optimizer (lines 3-4 in Alg. 1).
@@ IV Domain-Agnostic Co-Evolution of Parameterized Search (DACE)
DACE is a general-purpose, off-the-shelf approach for constructing PAPs for binary optimization problems. Same as CEPS [11], DACE evolves two competing populations: a configuration population (PAP) and a problem instance population. However, in DACE, problem instances are represented as NIRs, and are generated through the mutation operator described in Section III. Additionally, configurations in the PAP are identified from a configuration space defined by a parameterized algorithm. To eliminate the need for domain-specific parameterized algorithms, general-purpose EAs are used in DACE. Specifically, the biased random-key genetic algorithm (BRKGA) [38] is employed here, which has five configurable parameters: elite population size, offspring population size, mutant population size, elite bias for parent selection, and duplicate elimination flag. Details of these parameters are provided in Appendix B of the supplementary. The configuration space $\Theta$ consists of all possible combinations of these parameter values.
@@
Given the current PAP $P$, a so-called configuration mining process is iterated $n$ times (lines 10-15). In the $i$-th iteration, the $j$-th member algorithm is removed from $P$ to form $P^{\prime}$, where $j=i\mod K$ (lines 11-12). Then, an AAC procedure (SMAC [8] is used here, following CEPS) is employed to search for a new configuration $\theta^{\prime}$ within $\Theta$ that maximizes $\sum\nolimits_{m\in M}f(P^{\prime}\cup\{\theta^{\prime}\},m)$ (line 13), i.e., the performance of $P^{\prime}\cup\left\{\theta^{\prime}\right\}$ on the instance population $M$. After $n$ iterations, $n$ new configurations are obtained. In CEPS, the configuration whose inclusion yields the best-performing PAP is selected from these $n$ configurations to update $P$, and it is analogous to a mutation operation where the AAC procedure acts as a mutation operator that perturbs one member algorithm in $P$. DACE extends this approach. Specifically, a configuration set $\Psi$ is constructed by combining all $n$ new configurations with the $K$ configurations in $P$ (line 9 and line 14). DACE then examines all possible combinations of $K$ configurations from $\Psi$ and selects the combination with the best performance on $M$ to replace the current PAP (line 16). Since the performance of each configuration in $\Psi$ on each NIR in $M$ has already been evaluated, this enumeration does not require actually running the configurations on NIRs, but simply queries the previously recorded performance results. Given that $\left|\Psi\right|=K+n$, the total number of combinations to examine is a combinatorial value$\binom{n+K}{K}=\frac{\left(n+K\right)!}{K!n!}$, and this would take negligible time when $K$ and $n$ are not large (in our experiments, $K=4$ and $n=20$, with $\binom{n+K}{K}=10626$). As a result, this approach can potentially update all configurations in $P$ simultaneously, which is a more powerful mutation operation compared to CEPS’s single-configuration mutation. Since this improvement is not the primary focus of this work, the same PAP mutation mechanism is also applied to CEPS in our experiments to enhance its performance.
@@
The objective of evolving $M$ is to identify new NIRs that are challenging for the current PAP $P$. The fitness of each NIR $m$ is defined as $-f(P,m)$ (line 18). That is, NIRs on which $P$ performs poorly have higher fitness values. DACE begins by creating a copy $M^{\prime}$ of the instance population $M$ (line 19). Additionally, an empty set $M_{new}$ is initialized to store newly generated NIRs (line 20). DACE then repeats a so-called instance mining process $\left|M^{\prime}\right|/2$ times (lines 22-27). In each iteration, the mutation operator described in Alg. 1 is applied to an NIR $m$ randomly selected from $M^{\prime}$ to generate a new NIR $m^{\prime}$ (line 22). If no NIR in $M$ has a lower fitness than $m^{\prime}$, the mining process terminates (lines 24-25). Otherwise, an instance $m^{\star}$ with lower fitness than $m^{\prime}$ is randomly removed from $M$, and $m^{\prime}$ is added to $M_{new}$ (lines 26-27 ). When the mining process completes, the newly generated NIRs in $M_{new}$ are added to the instance population $M$ (line 29), which is used to obtain the new PAP in the next co-evolution round. The evolution of $M$ will be skipped in the last co-evolution round (line 17) because there is no need to generate more NIRs since the final PAP has been constructed completely.
@@ V Computational Studies
Extensive experiments are conducted on three binary optimization problems from real-world applications: the complementary influence maximization problem (CIMP) [39], compiler arguments optimization problem (CAOP) [21], and contamination control problem (CCP) [40]. Through the experiments, we aim to assess the potential of DACE by addressing the following research questions (RQs):
@@
For each problem class, problem instances were generated based on public benchmarks and divided into training and test sets. Following the experimental setup of CEPS [11], few-shot scenarios were simulated where the training set size is limited. Specifically, the training set for each problem class contained 5 instances, while the test set contained 100 instances with problem dimensions equal to or larger than training instances. Throughout the experiments, test instances were used solely to evaluate the generalization performance of PAPs obtained by DACE and the compared methods. The training instances were only used for PAP construction, regardless of which method used. The source code and benchmark instances for each problem class have been anonymously open-sourced at https://github.com/AnonymousSubmitBot/DACE.
@@ V-A Constraint Handling, Problem Classes and Instances
During the three problems, CIMP and CCP involve constraints. For these, constraint-handling techniques are applied. In CIMP, where the constraint limits the selection to at most $k$ seeds, solution repair [41] is employed – any solutions exceeding this limit are repaired by keeping only the first $k$ chosen seeds and discarding the rest. For CCP, established methods from prior studies [42] are adopted, integrating constraints as penalty components within the ojective function. Details of each problem class (including the objective function) and their problem instances are provided below.
@@ V-A1 Complementary Influence Maximization Problem (CIMP)
The rapid growth of online social platforms has made influence maximization an increasingly important problem. Given a social network $\mathcal{G}=\left(V,E,p\right)$, where $p$ specifies influence probabilities between nodes, the influence maximization problem (IMP) aims to find a seed set $S$ of $k$ nodes to maximize the expected number of active nodes. CIMP [39] extends IMP by introducing complementary users, making the influence analysis more complex due to collaborator interactions. Specifically, given a social network $\mathcal{G}$ and a complementary seed set $S_{A}$ for opinion $A$, CIMP aims to find a seed set $S_{B}$ of $k$ nodes from a candidate set $C\subset V$ with $\left|C\right|=d_{I}$ to maximize the spread of opinion $B$, making it a $d_{I}$-dimensional binary optimization problem. The interaction between opinions $A$ and $B$ is governed by parameters $q_{A|\emptyset}$, $q_{A|B}$, $q_{B|\emptyset}$, $q_{B|A}$, as detailed in [39].
In the binary solution representation $\boldsymbol{x}\in\{0,1\}^{d_{I}}$, $x_{i}=1$ indicates the $i$-th node is included into $S_{B}$, while $x_{i}=0$ indicates it is excluded. The objective function is to maximize the expected number of nodes activated by $S_{B}$, subject to selecting at most $k$ nodes:
$\begin{gathered}\max_{\boldsymbol{x}\in\left\{0,1\right\}^{d_{I}}}\textrm{ActiveNum}\left(\mathcal{G},S_{A},C,q_{A|\emptyset},q_{A|B},q_{B|\emptyset},q_{B|A},\boldsymbol{x}\right)\\ s.t.\quad\sum_{i=1}^{d_{I}}x_{i}\leq k\end{gathered}.$ (12)
Given a solution, if it exceeds $k$ nodes, it is first trimmed to $k$ nodes, and then Eq. (12) is evaluated through Monte Carlo simulations of the propagation process on the network.
For CIMP, problem instances were generated using three public social network benchmarks: Wiki [43], Facebook [44], and Epinions [43]. The Wiki benchmark was used for generating training instances, while test instances were generated from Epinions, Facebook and Wiki benchmarks, with the benchmark being randomly selected for each instance. For each $d_{I}$-dimensional CIMP instance, a candidate seed set $C$ of size $d_{I}$ was randomly selected from the node set $V$. The interaction parameters $\left\{q_{A|\emptyset},q_{A|B},q_{B|\emptyset},q_{B|A}\right\}$ were randomly selected from two default settings: $\left\{0.5,0.75,0.5,0.75\right\}$, and $\left\{0.5,0.25,0.5,0.25\right\}$. Finally, the training set contained five problem instances with $d_{I}=80$, and the test set contained 100 instances evenly split between $d_{I}=80$ and $d_{I}=100$.
@@ V-A2 Compiler Arguments Optimization Problem (CAOP)
CAOP [21] aims to minimize executable file size during software compilation, which is crucial in storage-limited environments. Software developers use compiler arguments to control various features that can reduce code size. The impact of these arguments varies depending on the source code being compiled. While appropriate arguments can effectively reduce file size, poor choices may lead to increased size. Therefore, selecting the right compiler arguments is critical for optimal results.
@@
For CAOP, problem instances were generated using the cbench and polybench-cpu [45] benchmarks, which contain 50 program source files in total. Among them 11 files were selected as training group and 26 files were selected as test group. To generate a training instance, a source file was randomly selected from the training group, while test instances were generated using random selections from the test group. Following [21], GCC was chosen as the compiler due to its widespread use. Among GCC’s 186 available arguments, $d_{I}$ arguments were randomly selected for each instance, with all other arguments being disabled. Finally, the training set contained five problem instances with $d_{I}=80$, and the test set contained 100 instances in total, evenly split between $d_{I}=80$ and $d_{I}=100$.
@@ V-A3 Contamination Control Problem (CCP)
CCP [40, 42] arises from the need for contamination prevention in the food production supply chain. Multiple stages are involved during food production, each of which can potentially introduce contamination. At stage $i$, taking mitigation measures can reduce contamination at a rate of random variable $\Gamma_{i}$, but incurs a cost of $c_{i}$. If no action is taken, the contamination rate will be $\alpha_{i}$. In a $d_{I}$-dimensional CCP instance, there are $d_{I}$ stages. For a binary solution $\boldsymbol{x}\in\{0,1\}^{d_{I}}$, $x_{i}=1$ indicates measures are taken at stage $i$ while $x_{i}=0$ representing no measures. Let $z_{i}$ denote the proportion of contaminated food at stage $i$, which is defined as $z_{i}=\alpha_{i}(1-x_{i})(1-z_{i-1})+(1-\Gamma_{i}x_{i})z_{i-1}$.
In CCP, the constraint limits the probability that contamination at each stage exceeds the upper limit $u_{i}$, approximated through $T$ Monte Carlo simulations. Following the previous work [42], this constraint is incorporated into the objective function as a penalty term $\frac{1}{T}\sum_{k=1}^{T}1_{\left\{z_{k}>u_{i}\right\}}$, and the final objective function is:
@@
In [42], CCP instances were generated with a dimension of 21, where $\lambda$ was selected from $\left\{0,10^{-4},10^{-2}\right\}$, while the random variables followed distributions: $\alpha\sim\textrm{Be}\left(1,\frac{17}{3}\right)$, $\Gamma\sim\textrm{Be}\left(1,\frac{7}{3}\right)$, and the initial contamination $z_{0}\sim\textrm{Be}\left(1,30\right)$. Since $\lambda$ was found to significantly influence instance characteristics, we adopted their instance generation approach and used $\lambda$ values to distinguish between training and test instances. Specifically, $\lambda$ was set to $10^{-4}$ for training instances and randomly selected from $\left\{0,10^{-2}\right\}$ for test instances. The training set consisted of five problem instances with $d_{I}=30$, while the test set contained 100 instances in total, evenly split between $d_{I}=30$ and $d_{I}=40$.
@@ V-B Compared Methods
To address RQ1, DACE was compared with several state-of-the-art PAP construction methods across all three problem classes. The main comparison was made with CEPS [11], which uses instance generation during PAP construction. The domain-specific instance mutation operator in CEPS was implemented exactly as the training instance generation mechanism described in Section V-A, thus generating in-distribution instances identical to the training set. This is to simulate that CEPS’s users possess comprehensive domain-specific knowledge of these three problem classes.
Additionally, GLOBAL [9] and PARHYDRA [18] were included as they represent state-of-the-art PAP construction that assumes sufficient training instances. PCIT [10] and CLUSTERING [17] were not included in the comparison due to their clustering mechanism being invalid with the limited training instances. To ensure fair comparison, all methods constructed PAPs based on the same configuration space defined by BRKGA, a general-purpose EA implemented using an open-source library [46]. This means that all constructed PAPs consist of BRKGA configurations. A manually constructed PAP (referred to as BRKGA-PAP) was also included as a baseline. This PAP contains four configurations: two recommended configurations from the open-source BRKGA library [46] and previous work [38], plus two additional configurations created by flipping the “eliminate_duplicates” parameter in the original two configurations.
To address RQ2, a SMARTEST-based PAP (referred to as SMARTEST-PAP) was included for comparison on CAOP. SMARTEST [21] is the state-of-the-art optimizer for CAOP, and its PAP variant is stronger than a single SMARTEST configuration. This PAP contains four SMARTEST configurations recommended in [21], each designed for instances with different characteristics. The specific configurations in BRKGA-PAP and SMARTEST-PAP are provided in Appendix C of the supplementary.

### v4

@@ 15     $\Psi\leftarrow\Psi\cup\left\{\theta^{\prime}\right\}$;
20  Assign the fitness of each $m\in M$ as f(P,m)-f(P,m);
@@ 22  $M_{new}\leftarrow\emptyset$;
25     Test $P$ on $m^{\prime}$ and assign the fitness of $m^{\prime}$ as $f(P,m^{\prime})$;
@@
28     $M\{$$\leftarrow M\backslash\left\{m^{\star}\right\}$;
@@ III-C NIR-based Instance Generation and Evaluation
Since the performance indicator $f$ typically lacks analytic forms and $\boldsymbol{e}^{\prime}$ is a real-valued vector, Eq. (10) represents a continuous black-box optimization problem. In this work, PGPE [beyer_evolution_2002], an evolution strategy (ES) method, is employed to optimize Eq. (10) (lines 1-2 in Alg. 1). Details of PGPE are provided in Appendix A of the supplementary. Note that other black-box continuous optimizers could also be used here, as the specific choice of optimizer is not central to our approach. Upon termination, the mutation operator returns the NIR $m^{\prime}$ specified by the best embedding $\boldsymbol{e}^{\prime}$ found by the optimizer (lines 3-4 in Alg. 1).
@@ IV Domain-Agnostic Co-Evolution of Parameterized Search (DACE)
DACE is a general-purpose, off-the-shelf approach for constructing PAPs for binary optimization problems. Same as CEPS [tang_few-shots_2021], DACE evolves two competing populations: a configuration population (PAP) and a problem instance population. However, in DACE, problem instances are represented as NIRs, and are generated through the mutation operator described in Section III. Additionally, configurations in the PAP are identified from a configuration space defined by a parameterized algorithm. To eliminate the need for domain-specific parameterized algorithms, general-purpose EAs are used in DACE. Specifically, the biased random-key genetic algorithm (BRKGA) [goncalves_biased_2011] is employed here, which has five configurable parameters: elite population size, offspring population size, mutant population size, elite bias for parent selection, and duplicate elimination flag. Details of these parameters are provided in Appendix B of the supplementary. The configuration space $\Theta$ consists of all possible combinations of these parameter values.
@@
[…]
@@
The objective of evolving $M$ is to identify new NIRs that are challenging for the current PAP $P$. The fitness of each NIR $m$ is defined as f(P,m)-f(P,m) (line 18). That is, NIRs on which $P$ performs poorly have higher fitness values. DACE begins by creating a copy $M^{\prime}$ of the instance population $M$ (line 19). Additionally, an empty set $M_{new}$ is initialized to store newly generated NIRs (line 20). DACE then repeats a so-called instance mining process $\left|M^{\prime}\right|/2$ times (lines 22-27). In each iteration, the mutation operator described in Alg. 1 is applied to an NIR $m$ randomly selected from $M^{\prime}$ to generate a new NIR $m^{\prime}$ (line 22). If no NIR in $M$ has lower quality than $m^{\prime}$, indicating that the generated instance is not challenging for the current PAP, the mining process terminates (lines 24–25). Otherwise, an instance $m^{\star}$ with lower fitness than $m^{\prime}$ is randomly removed from $M$, and $m^{\prime}$ is added to $M_{new}$ (lines 26-27 ). When the mining process completes, the newly generated NIRs in $M_{new}$ are added to the instance population $M$ (line 29), which is used to obtain the new PAP in the next co-evolution round. The evolution of $M$ will be skipped in the last co-evolution round (line 17) because there is no need to generate more NIRs since the final PAP has been constructed completely.
@@ V Computational Studies
Extensive experiments are conducted on three binary optimization problems from real-world applications: the complementary influence maximization problem (CIMP) [lu_competition_2015], compiler arguments optimization problem (CAOP) [jiang_smartest_2022], and contamination control problem (CCP) [hu_contamination_2010]. Through the experiments, we aim to assess the potential of DACE by addressing the following research questions (RQs):
@@
For each problem class, problem instances were generated based on public benchmarks and divided into training and test sets. Following the experimental setup of CEPS [tang_few-shots_2021], few-shot scenarios were simulated where the training set size is limited. Specifically, the training set for each problem class contained 5 instances, while the test set contained 100 instances with problem dimensions equal to or larger than training instances. Throughout the experiments, test instances were used solely to evaluate the generalization performance of PAPs obtained by DACE and the compared methods. The training instances were only used for PAP construction, regardless of which method used. The source code and benchmark instances for each problem class have been anonymously open-sourced at https://github.com/MetaronWang/DACE.
@@ V-A Constraint Handling, Problem Classes and Instances
During the three problems, CIMP and CCP involve constraints. As noted in Section III-B, NIR does not handle constraints directly. Instead, problem-specific constraint-handling techniques are applied here. Specifically, solution repair [coello2002theoretical] is employed to handle the constraint on the size of the selected seed set in CIMP. For CCP, following the previous study [oh_combinatorial_2019], the contamination probability constraint is converted to a penalty component. Details on each problem class, including their objective functions, constraint-handling techniques, and problem instances, are provided below.
@@ V-A1 Complementary Influence Maximization Problem (CIMP)
The rapid growth of online social platforms has made influence maximization an increasingly important problem. Given a social network $\mathcal{G}=\left(V,E,p\right)$, where $V$ denotes the nodes, $E$ denotes the edges, and $p$ specifies influence probabilities between nodes, the influence maximization problem (IMP) aims to find a seed set $S$ of $k$ nodes (each selected node is called a “seed”) to maximize the expected number of active nodes. CIMP [lu_competition_2015] extends IMP by introducing complementary users, making the influence analysis more complex due to collaborator interactions. Specifically, given a social network $\mathcal{G}$ and a complementary seed set $S_{A}$ for opinion $A$, CIMP aims to find a seed set $S_{B}$ of $k$ nodes from a candidate set $C\subset V$ with $\left|C\right|=d_{I}$ to maximize the spread of opinion $B$, making it a $d_{I}$-dimensional binary optimization problem. The interaction between opinions $A$ and $B$ is governed by parameters $q_{A|\emptyset}$, $q_{A|B}$, $q_{B|\emptyset}$, $q_{B|A}$, as detailed in [lu_competition_2015].
In the binary solution representation $\boldsymbol{x}\in\{0,1\}^{d_{I}}$, $x_{i}=1$ indicates the $i$-th node is included into $S_{B}$, while $x_{i}=0$ indicates it is excluded. The objective function is to maximize the expected number of nodes activated by $S_{B}$, subject to selecting at most $k$ seeds:
$\begin{gathered}\max_{\boldsymbol{x}\in\left\{0,1\right\}^{d_{I}}}\textrm{ActiveNum}\left(\mathcal{G},S_{A},C,q_{A|\emptyset},q_{A|B},q_{B|\emptyset},q_{B|A},\boldsymbol{x}\right)\hskip-10.0pt\\ s.t.\quad\Sigma_{i=1}^{d_{I}}x_{i}\leq k\end{gathered}$ (12)
Given a solution, if it exceeds $k$ selected seeds, it is repaired by keeping only the first $k$ chosen seeds and discarding the rest. Then, Eq. (12) is evaluated through Monte Carlo simulations of the propagation process on the network.
For CIMP, problem instances were generated using three public social network benchmarks: Wiki [leskovec2010predicting], Facebook [mcauley2012learning], and Epinions [leskovec2010predicting]. The Wiki benchmark was used for generating training instances, while test instances were generated from Epinions, Facebook and Wiki benchmarks, with the benchmark being randomly selected for each instance. For each $d_{I}$-dimensional CIMP instance, a candidate seed set $C$ of size $d_{I}$ was randomly selected from the node set $V$. The interaction parameters $\left\{q_{A|\emptyset},q_{A|B},q_{B|\emptyset},q_{B|A}\right\}$ were randomly selected from two default settings: $\left\{0.5,0.75,0.5,0.75\right\}$, and $\left\{0.5,0.25,0.5,0.25\right\}$. The maximum seed set size $k$ is an integer randomly selected from $\left[0.2d_{I},0.6d_{I}\right]$. Finally, the training set contained five problem instances with $d_{I}=80$, and the test set contained 100 instances evenly split between $d_{I}=80$ and $d_{I}=100$.
@@ V-A2 Compiler Arguments Optimization Problem (CAOP)
CAOP [jiang_smartest_2022] aims to minimize executable file size during software compilation, which is crucial in storage-limited environments. Software developers use compiler arguments to control various features that can reduce code size. The impact of these arguments varies depending on the source code being compiled. While appropriate arguments can effectively reduce file size, poor choices may lead to increased size. Therefore, selecting the right compiler arguments is critical for optimal results.
@@
For CAOP, problem instances were generated using the cbench and polybench-cpu [fursin2009collective] benchmarks, which contain 50 program source files in total. Among them 11 files were selected as training group and 26 files were selected as test group. To generate a training instance, a source file was randomly selected from the training group, while test instances were generated using random selections from the test group. Following [jiang_smartest_2022], GCC was chosen as the compiler due to its widespread use. Among GCC’s 186 available arguments, $d_{I}$ arguments were randomly selected for each instance, with all other arguments being disabled. Finally, the training set contained five problem instances with $d_{I}=80$, and the test set contained 100 instances in total, evenly split between $d_{I}=80$ and $d_{I}=100$.
@@ V-A3 Contamination Control Problem (CCP)
CCP [hu_contamination_2010, oh_combinatorial_2019] arises from the need for contamination prevention in the food production supply chain. Multiple stages are involved during food production, each of which can potentially introduce contamination. At stage $i$, taking mitigation measures can reduce contamination at a rate of random variable $\Gamma_{i}$, but incurs a cost of $c_{i}$. If no action is taken, the contamination rate will be $\alpha_{i}$. In a $d_{I}$-dimensional CCP instance, there are $d_{I}$ stages. For a binary solution $\boldsymbol{x}\in\{0,1\}^{d_{I}}$, $x_{i}=1$ indicates measures are taken at stage $i$ while $x_{i}=0$ representing no measures. Let $z_{i}$ denote the proportion of contaminated food at stage $i$, which is defined as $z_{i}=\alpha_{i}(1-x_{i})(1-z_{i-1})+(1-\Gamma_{i}x_{i})z_{i-1}$.
In CCP, the constraint limits the probability that contamination at each stage exceeds the upper limit $u_{i}$, approximated through $T$ Monte Carlo simulations. Following the previous work [oh_combinatorial_2019], this constraint is incorporated into the objective function as a penalty term $\frac{1}{T}\sum_{k=1}^{T}1_{\left\{z_{k}>u_{i}\right\}}$, and the final objective function is:
@@
In [oh_combinatorial_2019], CCP instances were generated with a dimension of 21, where $\lambda$ was selected from $\left\{0,10^{-4},10^{-2}\right\}$, while the random variables followed distributions: $\alpha\sim\textrm{Be}\left(1,\frac{17}{3}\right)$, $\Gamma\sim\textrm{Be}\left(1,\frac{7}{3}\right)$, and the initial contamination $z_{0}\sim\textrm{Be}\left(1,30\right)$. Since $\lambda$ was found to significantly influence instance characteristics, we adopted their instance generation approach and used $\lambda$ values to distinguish between training and test instances. Specifically, $\lambda$ was set to $10^{-4}$ for training instances and randomly selected from $\left\{0,10^{-2}\right\}$ for test instances. The training set consisted of five problem instances with $d_{I}=30$, while the test set contained 100 instances in total, evenly split between $d_{I}=30$ and $d_{I}=40$.
@@ V-B Compared Methods
To address RQ1, DACE was compared with several state-of-the-art PAP construction methods across all three problem classes. The main comparison was made with CEPS [tang_few-shots_2021], which uses instance generation during PAP construction. The domain-specific instance mutation operator in CEPS was implemented exactly as the training instance generation mechanism described in Section V-A, thus generating in-distribution instances identical to the training set. This is to simulate that CEPS’s users possess comprehensive domain-specific knowledge of these three problem classes.
Additionally, GLOBAL [lindauer_automatic_2017] and PARHYDRA [xu_hydra_2010] were included as they represent state-of-the-art PAP construction that assumes sufficient training instances. PCIT [liu_automatic_2019] and CLUSTERING [kadioglu_isac_2010] were not included in the comparison due to their clustering mechanism being invalid with the limited training instances. To ensure fair comparison, all methods constructed PAPs based on the same configuration space defined by BRKGA, a general-purpose EA implemented using an open-source library [blank_pymoo_2020]. This means that all constructed PAPs consist of BRKGA configurations. A manually constructed PAP (referred to as BRKGA-PAP) was also included as a baseline. This PAP contains four configurations: two recommended configurations from the open-source BRKGA library [blank_pymoo_2020] and previous work [goncalves_biased_2011], plus two additional configurations created by flipping the “eliminate_duplicates” parameter in the original two configurations.
To address RQ2, a SMARTEST-based PAP (referred to as SMARTEST-PAP) was included for comparison on CAOP. SMARTEST [jiang_smartest_2022] is the state-of-the-art optimizer for CAOP, and its PAP variant is stronger than a single SMARTEST configuration. This PAP contains four SMARTEST configurations recommended in [jiang_smartest_2022], each designed for instances with different characteristics. The specific configurations in BRKGA-PAP and SMARTEST-PAP are provided in Appendix C of the supplementary.